OpenStudy (anonymous):

Someone please help, Simplify 8 to the negative 1 third power

OpenStudy (anonymous):

do you mean \[(8^{-1})^{3}\] ??

OpenStudy (anonymous):

\[8^{-1/3}=(\frac{ 1 }{ 8 })^{1/3}=\sqrt[3]{1/8}=1/2\]

OpenStudy (anonymous):

8^(-1/3) x^-1 = 1/(x^1) so 8^(-1/3) = 1/(8^(1/3)) 8 = 2^3 so 1/((2^3)^(1/3)) equals 1/((x^3*(1/3))) equals 1/x^(3/3) equals 1/x^1 equals 1/x

OpenStudy (anonymous):

Perfect answer @yuendat99 ... it is actually .5

OpenStudy (anonymous):

Thanks guys. (:

OpenStudy (anonymous):

for some reason i substituted an x instead of my 2 so at the end it is 1/2

OpenStudy (anonymous):

Okay thanks